# VimTS: A Unified Video and Image Text Spotter for Enhancing the Cross-domain Generalization

### v3

@@ 4.1 Implementation Details
We pre-train the model on a combination of Curved SynthText [24], ICDAR-MLT [76], and the corresponding datasets with $240$K iterations. The base learning rate is $1\times 10^{-4}$ and reduced to $1\times 10^{-5}$ at the $180$K-th iteration and $1\times 10^{-6}$ at $210$K-th iteration. Then, the model is fine-tuned on the corresponding real datasets. We use $N=100$ as the maximum number of predictions. The max length of recognition queries is $25$. The weight for the classification loss $\alpha_{c}$ is $2.0$. The weight of the $\ell_{1}$ loss is $5.0$ and the GIoU loss is $2.0$. The polygon loss weight $\alpha_{p}$ and the recognition loss weight $\alpha_{r}$ are both set to $1.0$. The focal loss parameters $\alpha$ and $\gamma$ are $0.25$ and $2.0$, respectively. The number of both encoder and decoder layers is 6.
@@
Method Tuned Params Full-Tuning PQGM Adapter TA Adapter $TT_{W}$rightarrow $IC15$ $CTW_{L}$rightarrow $TT_{L}$ $TT_{W}$rightarrow $CTW_{W}$
Baseline - - - - - 65.0 50.5 80.9
Baseline+ 56.6M (100%) - - - 62.5 (-2.5) 36.6 (-13.9) 71.1 (-9.8)
Baseline+ 56.9M (100%) - - 67.6 (+2.6) 54.0 (+3.5) 81.5 (+0.6)
Baseline+ 0.7M (1.2%) - - 67.2 (+2.2) 53.9 (+3.4) 81.3 (+0.4)
Baseline+ 1.7M (3.0%) - - 68.0 (+3.0) 54.5 (+4.0) 82.2 (+1.3)
@@ 4.2 Comparison with State-of-the-arts Methods
Cross-Domain Text Spotting in Arbitrarily-Shaped Text. Following the previous method [7], we test our method on two arbitrarily-shaped text benchmarks (Total-Text and CTW1500) to verify the generalization ability of our method for arbitrarily-shaped cross-domain scene text spotting. For the text detection task, as shown in Tab. II. VimTS outperforms the previous methods with $91.5\%$ in terms of the H-mean metric on line-level Total-Text dataset, $1.9\%$ higher than the DeepSolo. The result of cross-domain text spotting is presented in Tab. III. Notably, even without line-level annotations from Total-Text for training, our method demonstrates promising performance in line-level Totaltext. Furthermore, our method outperforms the previous cross-domain method, FastTCM, by an average of $6.9\%$ in six cross-domain benchmarks. For text detection, our method outperforms the state-of-the-art approach by an average of $1.0\%$ across six cross-domain benchmarks. Similarly, in text spotting, our method demonstrates superiority over the state-of-the-art by an average of $2.6\%$ across the same benchmarks. These results demonstrate the strong generalization ability of our method.
4.3 Ablation Study in Cross-Domain Text Spotting
We conduct ablation studies in cross-domain text spotting using Total-Text-line, ICDAR2015, and CTW1500-word to evaluate the effectiveness of the proposed method. We directly train the model using both TotalText and CTW1500.
@@ 4.3 Ablation Study in Cross-Domain Text Spotting
TABLE V: Video text detection results in ICDAR2013 video dataset.
@@
TABLE VI: Video text detection results in DSText v2.
@@
TABLE VII: Zero-shot video text spotting on ICDAR2013 video. All methods use the same training set and evaluate on each frame of the video.
@@ 5 Experiments on Video Text Spotting
TABLE VIII: Video Text Spotting results. ‘M-Tracked’ and ‘M-Lost’ denote ‘Mostly Tracked’ and ‘Mostly Lost’, respectively.
@@ 5.1 Implementation Details
Following previous methods [39, 41], in our experiments, we develop a two-stage training manner. We first train our model using image-level datasets, which include ICDAR2013 [9], ICDAR2015 [8], Total-Text [10], and MLT2017 [76]. Consistent with prior works [39, 41], we incorporate random shift [91] to generate video clips with pseudo tracks in static images. For using the video data, we utilize the proposed synthetic video data. We jointly tune the PQGM and task-aware adapter on both image-level and video-level data building upon the pre-trained model on images. We utilize the AdamW [92] with a learning rate of $2\times 10^{-4}$ as the optimizer. Following previous work [41], we implement sequential data augmentation strategies, encompassing the following steps: 1) Random resizing of the image, with the shorter dimension chosen from a range of 608 to 992 pixels (in intervals of 32), while ensuring the longest dimension does not exceed 1536 pixels; 2) Random cropping, designed to preserve the entirety of the text without any truncation.
@@ 5.2 Zero-shot Transfer from Image to Video
To evaluate our method’s ability in video text spotting, we conduct a comparison with a previous image-level text spotting method using the ICDAR2013 video dataset [9] in a zero-shot manner. All methods are trained only using the image-level data and directly evaluated on evaluate on each frame of the ICDAR2013 video in a zero-shot manner. The results are shown in Tab. VII. VimTS achieves $53.6\%$ in the F-measure metrics, outperforming the previous image-level text spotting method by $8.2\%$ at least. These results demonstrate the superior generalization ability of our proposed method in the domain of video text spotting.
@@ 5.3 Comparison with State-of-the-arts Methods
Video text detection. For video text detection, similar to previous works [41, 42], we evaluate our method in the popular public dataset ICDAR2013 [9]. The results are presented in Tab. V. For the ICDAR2013 dataset, VimTS outperforms previous methods in all metrics. The proposed VimTS surpasses the previous end-to-end method TransDETR [41] by $2.3\%$ on the F-measure metric, leveraging solely image-level data and VTD-368k for training. Notably, VimTS outperforms previous methods even when trained solely on image-level data, highlighting the robust generalization and effectiveness of our method. Additionally, we also evaluate our method on the challenge dataset DSText v2 [90]. As shown in Tab. VI, our method achieves $64.2\%$ on the F-measure metric, which surpasses the previous state-of-the-art method by $2.1\%$. Additionally, our method achieves $63.7\%$ on the F-measure metric in a zero-shot manner, demonstrating the effectiveness of the proposed method.
Video text Spotting. For tasks of video text spotting, we evaluate our method on ICDAR2015 and DSText. The results are presented in Tab. VIII. In the scene text video benchmark ICDAR2015, VimTS surpasses the previous state-of-the-art method by $1.3\%$ in terms of the $ID_{F1}$ metrics, leveraging solely image-level data and VTD-368k for training. On the DSText v2 dataset, with a zero-shot manner, VimTS outperforms the previous end-to-end method TransDETR by $1.2\%$, $3.7\%$ in $ID_{F1}$, MOTP, respectively. Additionally, as shown in Tab. VIII, using real video data can further enhance the performance of our method. These results demonstrate the effectiveness of our method.
@@ 6 Discussion and Limitation
Scene text spotting has seen rapid development in recent years, yielding impressive results, particularly in uni-domain text spotting. However, cross-domain text spotting continues to present challenges, indicating areas for improvement. Our observation reveals that various models exhibit distinct behaviors when applied to cross-domain text spotting compared to their performance in uni-domain scenarios.
Recently, large multi-modal models have raised much attention for their robust generalization capabilities. To further show the effectiveness of our method, we conduct a cross-domain experiment on ICDAR2015 to compare with the large multi-modal models. The evaluation process is following  [93]. The results are presented in Tab. IX. The findings show the importance of developing scene text spotting methods for specific tasks. Such a specialized approach is not only more efficient in terms of requiring fewer parameters but also necessitates less training data compared to the broader application of Large Multimodal Models.
We also provide some failure cases in Fig. 10. We can find that even if the text instances in the previous frame have already been located and recognized. However, due to text instances moving quickly, leading to the text instances becoming significantly blurred, the model fails to locate and recognize them in the current frame. Therefore, the problem of blur of high-speed moving text is worth exploring in the future.
Figure 10: Failure cases. When the motion amplitude is too large, the text instances are blurred seriously. Even if the text instances are correctly located and recognized in the previous frame, the model fails to locate and recognize them in the current frame.
TABLE IX: Cross-domain text spotting on ICDAR2015 compared with the MLLMS. All results are tested on the ‘None’ lexicon. OCR-related data represents the training data associated with OCR.
@@ 7 Conclusion
In this paper, we introduce VimTS which enhances cross-domain text spotting performance by improving the synergy among hierarchical text spotting tasks, which include word-level, line-level, and video-level text spotting. VimTS enhances the generalization ability of the model by jointly optimizing different tasks across different scenarios. Extensive experiments on a wide range of cross-domain benchmarks, consistently demonstrate that our method outperforms previous state-of-the-art approaches by significant margins. It is worth mentioning that our method demonstrates it is viable that still text images can be learned to be well transferred to video text images. Since still images require significantly less annotation effort compared to video image, exploring methods to bridge the domain gaps will be highly valuable. Furthermore, we demonstrate that current Large Multimodal Models still face limitations in cross-domain text spotting. Using fewer parameters and less data to improve the generalization of Large Multimodal Models in text spotting is worth further exploration.
@@ Acknowledgement
This research was in part supported by National Key R&D Program of China (No.  2022ZD0118700), NSFC (No.  62225603, 62206104, 61936003), Zhuhai Industry Core and Key Technology Research Project (No.  2220004002350).

### v4

@@ 4.1 Implementation Details
We pre-train the model on a combination of Curved SynthText [24], ICDAR-MLT [80], and the corresponding datasets with $240$k iterations. The base learning rate is $1\times 10^{-4}$ and reduced to $1\times 10^{-5}$ at the $180$k-th iteration and $1\times 10^{-6}$ at $210$k-th iteration. Then, the model is fine-tuned on the corresponding real datasets. We use $N=100$ as the maximum number of predictions. The max length of recognition queries is $25$. The weight for the classification loss $\alpha_{c}$ is $2.0$. The weight of the $\ell_{1}$ loss is $5.0$ and the GIoU loss is $2.0$. The polygon loss weight $\alpha_{p}$ and the recognition loss weight $\alpha_{r}$ are both set to $1.0$. The focal loss parameters $\alpha$ and $\gamma$ are $0.25$ and $2.0$, respectively. The number of both encoder and decoder layers is 6. We obtain the TotalText-line and CTW1500-word by manually labeling, which has been made open-sourced.
@@
For the ‘Zero-shot’ setting, training requires approximately 45 hours on 8 RTX 3090 GPUs. The ‘with VTD-368k’ setting takes around 80 hours, and ‘with video data’ requires about 90 hours on the same setup.
@@
Method Tuned Params Full-Tuning PQGM Adapter TA Adapter $TT_{W}$rightarrow $IC15$ $CTW_{L}$rightarrow $TT_{L}$ $TT_{W}$rightarrow $CTW_{W}$ Training Time
Baseline - - - - - 65.0 50.5 80.9 -
Baseline+ 56.6M (100%) - - - 62.5 (-2.5) 36.6 (-13.9) 71.1 (-9.8) 210 min
Baseline+ 56.9M (100%) - - 67.6 (+2.6) 54.0 (+3.5) 81.5 (+0.6) 210 min
Baseline+ 0.7M (1.2%) - - 67.2 (+2.2) 53.9 (+3.4) 81.3 (+0.4) 140 min
Baseline+ 1.7M (3.0%) - - 68.0 (+3.0) 54.5 (+4.0) 82.2 (+1.3) 140 min
@@ 4.2 Comparison with State-of-the-arts Methods
Cross-Domain Text Spotting in Arbitrarily-Shaped Text. Following the previous method [7], we test our method on two arbitrarily-shaped text benchmarks (TotalText and CTW1500) to verify the generalization ability of our method for arbitrarily-shaped cross-domain scene text spotting. For the text detection task, as shown in Tab. II. VimTS outperforms the previous methods with $91.5\%$ in terms of the H-mean metric on line-level TotalText dataset, $1.9\%$ higher than the DeepSolo. The result of cross-domain text spotting is presented in Tab. III. Notably, even without line-level annotations from TotalText for training, our method demonstrates promising performance in line-level Totaltext. Furthermore, our method outperforms the previous cross-domain method, FastTCM, by an average of $6.9\%$ in six cross-domain benchmarks. For text detection, our method outperforms the state-of-the-art approach by an average of $1.0\%$ across six cross-domain benchmarks. Similarly, in text spotting, our method demonstrates superiority over the state-of-the-art by an average of $2.6\%$ across the same benchmarks. These results demonstrate the strong generalization ability of our method.
4.3 Ablation Study
We conduct ablation studies in cross-domain text spotting using TotalText-line, ICDAR2015, and CTW1500-word to evaluate the effectiveness of the proposed method. We directly train the model using both TotalText and CTW1500.
@@ 4.3 Ablation Study
Additionally, it’s noteworthy that fine-tuning the model with the respective real datasets typically takes about 3.5 hours. However, using PQGM with the task-aware adapter can further reduce this fine-tuning time to approximately 140 minutes, as shown in Tab. IV.
@@
Effectiveness of the order of detection and recognition queries. To further validate the influence of the order, we conduct an experiment as shown in Tab. V. The results show that detection before recognition is slightly better than recognition first setting.
TABLE V: Ablation study of the order of aggregating information in task-aware adapter.
Method Det. First $TT_{W}$-$IC15$ $CTW_{L}$-$TT_{L}$ $TT_{W}$-$CTW_{W}$
Method - 67.5 54.1 81.9
Method 68.0 54.5 82.2
TABLE VI: Ablation study of using different training data. ‘M-Tracked’ and ‘M-Lost’ denote ‘Mostly Tracked’ and ‘Mostly Lost’, respectively.
Data End to End Video Text Spotting/%
VTD-368k 53.7 30.9 74.6 29.1 47.2
image-level data 73.2 61.0 77.1 52.2 25.7
image-level data + VTD-368k 74.1 61.8 77.3 52.7 25.6
image-level data + ICDAR2015 74.5 62.0 77.5 52.6 25.2
image-level data + ICDAR2015 + VTD-368k 75.0 62.3 77.8 52.9 25.3
TABLE VII: Video text detection results in ICDAR2013 video dataset. The video data represent the training dataset of ICDAR2015 video.
@@
TABLE VIII: Video text detection results in DSText v2. The video data represents the training dataset of DSText v2.
@@
TABLE IX: Zero-shot video text spotting on ICDAR2013 video. All methods use the same training set and evaluate on each frame of the video.
@@ 5 Experiments on Video Text Spotting
TABLE X: Video Text Spotting results. The video data represent the corresponding training data of the benchmarks.
@@ 5.1 Implementation Details
Following previous methods [37, 39], in our experiments, we develop a two-stage training manner. We first train our model using image-level datasets, which include ICDAR2013 [9], ICDAR2015 [8], TotalText [10], and MLT2017 [80]. Consistent with prior works [37, 39], we incorporate random shift [95] to generate video clips with pseudo tracks in static images. For using the video data, we utilize the proposed synthetic video data. We jointly tune the PQGM and task-aware adapter on both image-level and video-level data building upon the pre-trained model on image text spotting. We utilize the AdamW [96] with a learning rate of $2\times 10^{-4}$ as the optimizer. Following previous work [39], we implement sequential data augmentation strategies, encompassing the following steps: 1) Random resizing of the image, with the shorter dimension chosen from a range of 608 to 992 pixels (in intervals of 32), while ensuring the longest dimension does not exceed 1536 pixels; 2) Random cropping, designed to preserve the entirety of the text without any truncation.
@@ 5.2 Zero-shot Transfer from Image to Video
To evaluate our method’s ability in video text spotting, we conduct a comparison with a previous image-level text spotting method using the ICDAR2013 video dataset [9] in a zero-shot manner. All methods are trained only using the image-level data and directly evaluated on each frame of the ICDAR2013 video in a zero-shot manner. The results are shown in Tab. IX. VimTS achieves $53.6\%$ in the F-measure metrics, outperforming the previous image-level text spotting method by $8.2\%$ at least. These results demonstrate the superior generalization ability of our proposed method in the domain of video text spotting.
@@ 5.3 Comparison with State-of-the-arts Methods
Video text detection. For video text detection, similar to previous works [39, 40], we evaluate our method in the popular public dataset ICDAR2013 [9]. The results are presented in Tab. VII. For the ICDAR2013 dataset, VimTS outperforms previous methods in all metrics. The proposed VimTS surpasses the previous end-to-end method TransDETR [39] by $2.3\%$ on the F-measure metric, leveraging solely image-level data and VTD-368k for training. Notably, VimTS outperforms previous methods even when trained solely on image-level data, highlighting the robust generalization and effectiveness of our method. Additionally, we also evaluate our method on the challenge dataset DSText v2 [94]. As shown in Tab. VIII, our method achieves $64.2\%$ on the F-measure metric, which surpasses the previous state-of-the-art method by $2.1\%$. Additionally, our method achieves $63.7\%$ on the F-measure metric in a zero-shot manner.
Video text Spotting. For task of video text spotting, we evaluate our method on ICDAR2015 and DSText v2. The results are presented in Tab. X. In the scene text video benchmark ICDAR2015, VimTS surpasses the previous state-of-the-art method by $1.3\%$ in terms of the $ID_{F1}$ metrics, leveraging solely image-level data and VTD-368k for training. On the DSText v2 dataset, with a zero-shot manner, VimTS outperforms the previous end-to-end method TransDETR by $1.2\%$, $3.7\%$ in $ID_{F1}$, MOTP, respectively. Additionally, as shown in Tab. X, using real video data can further enhance the performance of our method.
@@
5.4 Effectiveness of VTD-368k
Validating VTD-368k without real data. To further evaluate the effectiveness of VTD-368k. We conduct experiments with different dataset settings on ICDAR 2015 dataset, as shown in Tab. VI. For scenarios lacking real text images and text videos, we can train the model using only our synthetic data. It can be seen that even when using only synthetic data, the model is still able to perform video text spotting. The results indicate that synthetic data presents acceptable performance in scenarios where real data is not available.
Improving compactness by using VTD-368k. We conducted a series of experiments on the DSText v2 dataset, employing various IoU thresholds. The ground truth and evaluation criteria were sourced from the official website. The experimental results are summarized in Tab. XI. Our observations reveal that, in the absence of VTD-368k, there is a significant decline in performance as the evaluation criteria become more stringent, specifically when the IoU threshold increases from 0.5 to 0.7. Conversely, with the inclusion of VTD-368k, the model exhibits a much smaller performance degradation under the same conditions of increasing stringency in evaluation standards. These findings suggest that VTD-368k may be able to contribute to a greater quantity of accurate output from the model.
TABLE XI: Ablation study of using different training data on DSText v2.
Data IOU End to End Video Text Spotting on DSText v2/%
image-level data 0.5 12.5 -1.0 82.1 744 11362
image-level data + VTD-368k 0.5 13.2 -0.5 82.5 783 11031
image-level data 0.7 10.2 -14.4 78.1 246 11142
image-level data + VTD-368k 0.7 12.3 -1.3 81.9 758 10317
Different baselines of using VTD-368k. We conducted experiments to evaluate the performance of VTD-368k in comparison to other baselines. Specifically, we used TransDETR and TransVTSpotter as baseline models. The results, presented in Table XII, show that VTD-368k can improve the performance over both baselines.
TABLE XII: Ablation study of using the VTD-368k on TransDETR and TransVTSpotter (our reproduce).
Data End to End Video Text Spotting/%
TransDETR [39] 72.8 60.9 74.6 33.6 20.8
+ VTD-368k 74.0 61.8 75.3 36.3 19.2
TransVTSpotter [37] 61.1 45.8 73.6 29.2 37.2
+ VTD-368k 62.5 46.5 74.2 30.7 36.5
@@ 5.5 Visualization and Analysis
To validate whether different prompt queries can learn the task-specific features, we visualize the attention map of different prompt queries on image features. When word-level queries are used, the attention map focuses on word-level features, e.g., the model treats the features of the words ‘TOWNE’ and ‘MANOR’ separately, as presented in Fig. 10 (b). When line-level queries are used, the model combines the features of the words ‘TOWNE’ and ‘MANOR’ together, as shown in Fig. 10 (c). Intuitively, we can observe the difference between the two different queries.
Figure 10: The attention map of word-level and line-level prompt queries on image features.
@@ 6 Discussion and Limitation
Recently, Large Multi-modal Models (LMMs) have raised much attention for their robust generalization capabilities. To further show the effectiveness of our method, we conduct a cross-domain experiment on ICDAR2015 to compare with the LMMs. The evaluation process is following  [97]. The results are presented in Tab. XIII. The findings show the importance of developing scene text spotting methods for specific tasks. Such a specialized approach is not only more efficient in terms of requiring much fewer parameters but also necessitates significantly less training data compared to the broader application of LMMs.
We provide some failure cases of the VimTS in Fig. 12. Even when text instances in the previous frame have been successfully located and recognized, rapid movement can cause significant blurring, preventing the model from accurately locating and recognizing them in the current frame. Therefore, addressing the issue of blurred text in high-speed motion is an important direction for future exploration.
We also provide some failure synthetic samples of VTD-368k, as shown in Fig. 11. Errors can be observed in two scenes: cluttered scene and extreme motions scene. For the cluttered scene, due to the chaotic background, we are unable to composite the text into the video. For the extreme motions scene, the flipping or rapid movements make it difficult to establish the relationship between texts in consecutive frames during synthesis.
Figure 11: Failure cases of VTD-368k. (a) represents the cluttered-scene error. (b) represents the extreme motions error.
Figure 12: Failure cases of VimTS. When the motion amplitude is too large, the text instances are blurred seriously. Even if the text instances are correctly located and recognized in the previous frame, the model fails to locate and recognize them in the current frame.
TABLE XIII: Cross-domain text spotting on ICDAR2015 compared with the MLLMS. All results are tested on the ‘None’ lexicon. OCR-related data represents the training data associated with OCR.
@@ 7 Conclusion
In this paper, we present VimTS, which improves cross-domain text spotting by enhancing the synergy among hierarchical tasks, including word-level, line-level, and video-level spotting. VimTS boosts model generalization by jointly optimizing tasks across diverse scenarios. Extensive experiments on multiple cross-domain benchmarks consistently show that our method outperforms state-of-the-art approaches by significant margins. Notably, we demonstrate that text spotters trained on still images can be effectively transferred to video text spotting, offering a promising solution given the lower annotation effort required for still images. Additionally, we highlight that current large multimodal models still face challenges in cross-domain text spotting, and addressing blurred text in high-speed motion remains a crucial area for future research.
@@ Acknowledgement
This research was in part supported by National Key R&D Program of China (No.  2022ZD0118700, 2022YFC2305100), NSFC (No.  62225603, 62206104, 61936003), Zhuhai Industry Core and Key Technology Research Project (No.  2220004002350).